3. Your hospital board intends to acquire an equipment that will cost it Php 50,000,000. The board is thinking of financing this acquisition via by 70% internally generated funds and 30% loan at a fixed rate of 7% per annum. The hospital has consistently given out quarterly dividends of 4%. Calculate the hospitals weighted average cost of capital. 4. Assume the hospital has the following cash flows resulting from the machine Year acquisition: (0)70,000,000:(1)10,000,000:(2)30,000,000;(3)40,000,000:(4)25,000,000;(5)25,0 00.000; Calculate the projects NPV, Profitability Index and Discounted Payback Period.
Answer #1
3) Cost of equity = 4%*4 = 16.00%
WACC = 7%*30%+16%*70% = 13.30%
4) Year Cash Flow PVIF at 13.3% PV at 13.3% Cumulative PV
0 -70000000 1.00000 -70000000 -70000000
1 10000000 0.88261 8826125 -61173875
2 30000000 0.77900 23370146.51 -37803728
3 40000000 0.68756 27502378.94 -10301349
4 25000000 0.60685 15171215.22 4869866
5 25000000 0.53561 13390304.69 18260171
NPV = $18,260,171 18260170.69
PI = PV of cash inflows/Initial investment = 88260171/70000000 = 1.26
Discounted payback in years = 3+10301349/15171215 = 3.68
ALTERNATIVELY:
3) Cost of equity with quarterly compounding = 1.04^4-1 = 16.99%
WACC = 7%*30%+16.99%*70% = 13.99%
4) Year Cash Flow PVIF at 13.99% PV at 13.99% Cumulative PV
0 -70000000 1.00000 -70000000 -70000000
1 10000000 0.87727 8772699 -61227301
2 30000000 0.76960 23088076 -38139224
3 40000000 0.67515 27005967 -11133258
4 25000000 0.59229 14807202 3673944
5 25000000 0.51960 12989913 16663857
NPV = $16,663,857 16663857
PI = PV of cash inflows/Initial investment = 86663857/70000000 = 1.24
Discounted payback in years = 3+11133258/14807202 = 3.75